find the value of tan(alpha+beta), given that cot alpha=1/2, alpha€(pi,3pie/2), sec beta=-5/3, beta€(pie/2,pie). Also state the quadrant in which alpha+beta terminates.

Grade:11

1 Answers

Meet
6 years ago
cot theta is positive in 3rd quadrant and tan beta is neagtive in 2nd quadrant so therefore tan alpha=1/2 and tan beta=-4/3Then after putting values in the formula u will get tan(alpha+beta)=2/11